If the modom Variable Y denotes an individuals income, Paretos law claims that P(Yay)-(W) wherek is the entire populations
Answer #1

sample values in of random fy (Y) = oko ( 4 ) et ; y 7,K ,871 likelihood function of random sam 9, 92, ..., Yn is given as LI2 log L (8,0) = 0 2 n +nlogk - logp = 0 = logp - nlogk = logo = 1oo loga 1 = loga- logo (1.) gog ten : lôme = mLE of o = MLE-041 -ot y-of 900 => ELY) = OK0. ( y-04) - OK (O K ) (2 071) list - Ely) = OKO KI-O -(1-0) = -178 Now, for known k, to get mmlikelihood would whood would be marimum value of a lifo is known ). for the monimum le. L(koly, ...gr (K, 019,--yo) - (OKO)